# Optimized Lattice-Structured Flexible EIT Sensor for Tactile Reconstruction and Classification

### v1

@@ Abstract
Flexible electrical impedance tomography (EIT) offers a promising alternative to traditional tactile sensing approaches, enabling low-cost, scalable, and deformable sensor designs. Here, we propose an optimized lattice-structured flexible EIT tactile sensor incorporating a hydrogel-based conductive layer, systematically designed through three-dimensional coupling field simulations to optimize structural parameters for enhanced sensitivity and robustness. By tuning the lattice channel width and conductive layer thickness, we achieve significant improvements in tactile reconstruction quality and classification performance. Experimental results demonstrate high-quality tactile reconstruction with correlation coefficients up to 0.9275, peak signal-to-noise ratios reaching 29.0303 dB, and structural similarity indexes up to 0.9660, while maintaining low relative errors down to 0.3798. Furthermore, the optimized sensor accurately classifies 12 distinct tactile stimuli with an accuracy reaching 99.6%. These results highlight the potential of simulation-guided structural optimization for advancing flexible EIT-based tactile sensors toward practical applications in wearable systems, robotics, and human-machine interfaces.
@@
Electrical impedance tomography (EIT), lattice structure, tactile reconstruction, tactile classification
@@ I Introduction
Electronic skin (e-skin) is an emerging sensing technology that employs flexible and stretchable materials integrated with embedded sensors to replicate the sensory functions of human skin[1]. By enabling robots to interact with and accurately perceive their surroundings, even in complex and unstructured environments [2], e-skin represents a promising pathway toward embodied intelligence. Therefore, it has attracted significant attention for applications in human-machine interfaces (HMI) [3], health monitoring [4], and wearable devices [5]. Typically, e-skin systems consist of a flexible substrate integrated with various sensors capable of detecting stimuli such as pressure, temperature, and humidity. Among these, tactile sensors are particularly crucial for identifying contact locations and quantifying applied forces. Existing tactile sensors are typically based on piezoresistive[6], capacitive[7], optical[8], and magnetic[9] mechanisms. However, these approaches generally utilize arrays of discrete sensing units [10] that require intricate wiring and complex structures, leading to higher production costs, limited flexibility and scalability. Scaling such sensors to larger areas presents challenges such as reduced spatial resolution, higher costs and hardware complexity [11], thereby restricting their practical deployment in real-world applications.
@@
Conventional EIT-based tactile sensors usually adopt uniformly distributed conductive pathways, resulting in only small potential changes at the boundary electrodes when localised pressure is applied. This inherent characteristic limits the sensitivity of EIT-based tactile sensing systems. Recent research has explored novel structures to improve sensitivity, such as porous[17] and multi-layer structures[18]. However, these approaches often face manufacturing challenges that constrain scalability and durability. Furthermore, the lattice structure facilitates higher current density, enabling more pronounced and rapid conductivity changes under press deformation, thereby significantly improving the responsiveness of the sensing system [19]. However, current lattice-structured EIT tactile sensors face two key limitations[20]. First, critical lattice parameters, such as channel width and conductive layer thickness, are often selected empirically, without systematic optimization. Second, the influence of lattice structure on the EIT signal response has not been quantitatively analyzed or benchmarked against non-lattice configurations. These gaps limit the development of high-sensitivity, robust, and reproducible tactile sensors.
@@
  • We develop a flexible lattice-patterned EIT-based tactile sensor fabricated with hydrogel and silicone, whose lattice design parameters (specifically channel width and conductive layer thickness) are optimised through 3D-CFS to maximize sensing performance.
@@ II Principle of EIT-based tactile sensing
The principle of EIT-based tactile sensing is illustrated in Fig. 1. When localized pressure is applied to a sample substrate, it induces changes in conductivity within the touched regions, which in turn alter the electric potentials measured at the boundary electrodes. The objective of EIT-reconstruction in tactile sensing is to estimate the pressure-induced conductivity changes $\boldsymbol{\sigma}$ in the sensing region $\boldsymbol{\Omega}$ based on the voltage measurements $\boldsymbol{V}$. EIT involves solving both a forward and an inverse problem [24]. The forward problem predicts surface potentials for a given conductivity and applied current. This relationship is governed by:
@@
III Lattice-patterned Flexible Tactile Sensor
III-A 3D-CFS
@@ III Lattice-patterned Flexible Tactile Sensor
As shown in Fig. 2(a), the simulated sensor model is designed to be 100 × 100 × 9 mm and consists of three layers and 16 evenly distributed boundary electrodes. The substrate and sealing layers are non-conductive, and the sealing layer has a fixed thickness of 2 mm. The thickness of the substrate layer is changed with the intermediate conductive lattice layer to maintain a constant total model thickness of 9 mm. To evaluate and optimise sensor sensitivity, six different lattice channel widths and five different conductivity layer thicknesses were simulated to assess the influence of lattice channel widths and conductive layer thickness on sensor sensitivity.
As shown in Fig. 2(b), three different touch conditions are applied by varying the number of touch points from one to three. In each condition, 100 sets of touch phantoms are randomly generated, with press depths from 1 mm to 5 mm and touch radius from 6.25 mm to 8.75 mm.
@@
Different lattice parameters were calculated, with the conductive layer thickness from 1 mm to 5 mm and the channel width from 0 mm (no lattice pattern) to 6 mm. The result is as shown in Fig. 2d, each average value represents the average relative voltage change from 300 random touch phantoms across three touch conditions. The results demonstrate that a conductive layer thickness of 3 mm and a channel width of 4 mm consistently provided the highest sensitivity.
III-B Sensor Fabrication
Figure 4: Hydrogel fabrication process. (a) The solution contained 21.5 weight percent (wt%) lithium chloride (Sigma-Aldrich), 8 wtacrylamide (Sigma-Aldrich), 32.27 wtdeionised water and 37 wtethylene glycol (Sigma-Aldrich) was prepared. Ethylene glycol maintains hydrogel pliability across a wide temperature range, essential for stable sensor performance. (b) The solution was stirred at 60°C and 700 rpm for two hours to achieve uniform mixing and dissolution. (c) To initiate a polyacrylamide-based network, 1 wtpolyacrylamide (Sigma-Aldrich) was introduced, providing mechanical stability and enhancing sensor durability. (d) The solution was stirred overnight at 60 ${}^{\circ}\mathrm{C}$ and 700 rpm to ensure complete dissolution and formation of a uniform polymer network. (e) A crosslinking agent, 0.15 wtN,N’-methylenebisacrylamide (Sigma-Aldrich), was added to control the mechanical strength of the gel. Additionally, 0.08 wt2-hydroxy-2-methylpropiophenone (Sigma-Aldrich), a photoinitiator, was added to enable UV curing of the hydrogel. (f) The mixture was shielded from light and stirred for an additional two hours at 60 ${}^{\circ}\mathrm{C}$ and 700 rpm to ensure complete homogenization.
@@ III-B Sensor Fabrication
Based on the optimal design parameters identified through the 3D-CFS simulations, a conductive layer thickness of 3 mm and a lattice channel width of 4 mm, we proceeded to fabricate the lattice-structured EIT tactile sensor accordingly. The sensing layer is fabricated from a hydrogel primarily composed of acrylamide, a hydrophilic monomer known for polyacrylamide-based hydrogels, which retains water and ensures sensor functionality in humid conditions. The inclusion of lithium chloride, a common ionic conductor, further enhances the hydrogel’s ionic conductivity. The hydrogel precursor synthesis procedure is shown in Fig. 4.
As shown in Fig. 5, the hydrogel and silicone layers were integrated to fabricate the complete tactile sensor (Fig. 5h). To enhance durability and maintain long-term functionality, the entire structure was encapsulated with a silicone layer to prevent moisture loss from the hydrogel.
III-C Real-world Sensitivity Characterization and Validation
@@
To verify the optimized sensor parameters identified through CFS, we fabricated three sensors using the same procedures (refer to Fig. 5) and a consistent conductive layer thickness of 3 mm: two lattice-patterned sensors with different lattice parameters and one without the lattice structure for comparison. These sensors were subjected to a continuously increasing localized pressure with strains ranging from 0 mm to 4 mm, as illustrated in Fig. 6a. The pressure was applied through a circular indenter with a radius of 8.75 mm, simulating the typical contact area of a human fingertip. The pressing location, marked by the blue dot in the inset of Fig. 6a, remained consistent across all measurements. The results demonstrate that the sensor with optimised lattice parameters (channel width of 4 mm) consistently shows superior sensitivity compared to the non-lattice structured sensor and lattice-structure sensor with non-optimal parameters (channel width of 6 mm), aligning well with the simulation results. Specifically, at the press depth of 4 mm, the mean relative voltage change for the sensor with the optimized lattice (channel width of 4 mm) reaches approximately 0.0692, compared to  0.0613 for the 6 mm lattice sensor and only  0.0375 for the non-lattice sensor. This corresponds to an 84.5% improvement over the non-lattice structure and a 12.9% enhancement compared to the non-optimal parameters (channel width of 6 mm), clearly confirming the effectiveness of the optimized lattice pattern in enhancing sensitivity.
Additionally, tactile reconstruction results comparing the optimised lattice-patterned sensor and the sensor without the lattice channel are illustrated in Fig. 6b. The reconstructed images show that the optimised lattice-patterned sensor successfully reconstructs tactile information even at lower compressive strains, whereas the sensor without the lattice structure fails to identify tactile input at these shallow press depths. Furthermore, as the compressive strain increases, the lattice-patterned sensor consistently demonstrates superior reconstructions compared to the sensor without the lattice structure. These results further validate that the lattice-patterned sensor possesses higher sensitivity and enhanced tactile reconstruction performance.
III-D Cycling Test
@@
IV Tactile reconstruction
@@ IV-A Dataset and Network Training
To generate the dataset for the lattice-structured EIT sensor, we constructed a finite element model (FEM) using COMSOL Multiphysics and Matlab. We divide the sensing region with a size of 100 $\times$ 100 mm2 into 2304 elements (48 $\times$ 48) and set the background conductivity $\boldsymbol{\sigma_{0}}$ to 0.00312 S/m. Tactile interactions were simulated using circular regions with diameters ranging from 7.5 mm to 27.5 mm, varying in both location and intensity. The conductivity within these touch regions is randomly assigned values between $0.05\times\sigma_{0}$ and $2\times\sigma_{0}$. The complete dataset comprises 50,000 samples, divided into five subsets containing 10,000 samples each. These subsets correspond to different numbers of touch regions, ranging from 1 to 5 touch areas per sample. Each subgroup is further partitioned into training and validation sets using a 7:3 ratio. To enrich the training dataset by supplementing uncollected location information, we used a data augmentation approach to convert a single EIT measurement into 8 distinct readouts [26]. To further improve generalisation, additive Gaussian noise with Signal-to-Noise Ratios (SNRs) of 35, 40, 45, 50, 55, 60 and 65 dB was added to the seven augmented measurements, respectively. Finally, this process yielded a total of 280,000 samples for training and 15,000 samples for validation.
@@ IV-B Reconstruction Results and Discussion
Figure 7: Tactile reconstruction using experiment data. Images are normalised.
Figure 8: Tactile reconstruction of unseen real-world complex touches. Images are normalised.
As shown in Fig. 7, from Phantom 1 to Phantom 8, we applied 1 to 5 contacts at different positions using a linear robotic platform (DLE-RG-0001). The reconstruction results demonstrate that the sensor with the DNN model successfully reconstructs various touch patterns with good accuracy. To quantitatively assess reconstruction quality, we evaluate the results using correlation coefficients (CC), peak signal-to-noise ratio (PSNR), relative image error (RE), and structural similarity indexes (SSIM). Across all phantoms, the model achieves strong performance, with CC values reaching up to 0.8856, RE values as low as 0.4028, SSIM values up to 0.9648, and PSNR values reaching 29.0303 dB. These results indicate that the proposed sensor and model combination delivers high-fidelity reconstructions with strong consistency and visual clarity.
Beyond the above benchmark scenarios, we further validated the system’s performance on real-world measurements involving more complex and natural touch patterns. These include multi-finger contact, rectangular contact regions, and annular touch profiles. As shown in Fig. 8, the DNN model effectively reconstructs these diverse tactile configurations. Supplementary Video 1 further demonstrates the model’s real-time tactile reconstruction capabilities on selected phantoms, confirming its suitability for practical, interactive scenarios.
IV-C HMI Application

### v2

@@ Abstract
Flexible electrical impedance tomography (EIT) offers a promising alternative to traditional tactile sensing approaches, enabling low-cost, scalable, and deformable sensor designs. Here, we propose an optimized lattice-structured flexible EIT tactile sensor incorporating a hydrogel-based conductive layer, systematically designed through three-dimensional coupling field simulations to optimize structural parameters for enhanced sensitivity and robustness. By tuning the lattice channel width and conductive layer thickness, we achieve significant improvements in tactile reconstruction quality and classification performance. Experimental results demonstrate high-quality tactile reconstruction with correlation coefficients up to 0.9275, peak signal-to-noise ratios reaching 29.0303 dB, and structural similarity indexes up to 0.9660, while maintaining low relative errors down to 0.3798. Furthermore, the optimized sensor accurately classifies 12 distinct tactile stimuli with an accuracy reaching 99.6%. These results highlight the potential of simulation-guided structural optimization for advancing flexible EIT-based tactile sensors toward practical applications in wearable systems, robotics, and human-machine interfaces. All data are publicly available in Edinburgh DataShare with the identifier https://doi.org/10.7488/ds/7982.
@@
Electrical impedance tomography (EIT), tactile sensor, lattice structure, tactile reconstruction, tactile classification
@@ I Introduction
EElectronic skin (e-skin) is an emerging sensing technology that employs flexible and stretchable materials integrated with embedded sensors to replicate the sensory functions of human skin[1]. By enabling robots to interact with and accurately perceive their surroundings, even in complex and unstructured environments [2], e-skin represents a promising pathway toward embodied intelligence. Therefore, it has attracted significant attention for applications in human-machine interfaces (HMI) [3], health monitoring [4], and wearable devices [5]. Typically, e-skin systems consist of a flexible substrate integrated with various sensors capable of detecting stimuli such as pressure, temperature, and humidity. Among these, tactile sensors are particularly crucial for identifying contact locations and quantifying applied forces. Existing tactile sensors are typically based on piezoresistive[6], capacitive[7], optical[8], and magnetic[9] mechanisms. However, these approaches generally utilize arrays of discrete sensing units [10] that require intricate wiring and complex structures, leading to higher production costs, limited flexibility and scalability. Scaling such sensors to larger areas presents challenges such as reduced spatial resolution, higher costs and hardware complexity [11], thereby restricting their practical deployment in real-world applications.
@@
Conventional EIT-based tactile sensors usually adopt uniformly distributed conductive pathways, resulting in only small potential changes at the boundary electrodes when localised pressure is applied. This inherent characteristic limits the sensitivity of EIT-based tactile sensing systems. Recent research has explored novel structures to improve sensitivity, such as porous[17] and multi-layer structures[18]. However, these approaches often face manufacturing challenges that constrain scalability and durability. Furthermore, the lattice structure facilitates higher current density, enabling more pronounced and rapid conductivity changes under press deformation, thereby significantly improving the responsiveness of the sensing system [19]. However, current lattice-structured EIT tactile sensors face two key limitationsḞirst, critical lattice parameters, such as channel width and conductive layer thickness, are often selected empirically, without systematic optimization. Second, the influence of lattice structure on the EIT signal response has not been quantitatively analyzed or benchmarked against non-lattice configurations. These gaps limit the development of high-sensitivity, robust, and reproducible tactile sensors.
@@
  • We develop a flexible lattice-patterned EIT-based tactile sensor fabricated with hydrogel and silicone, whose lattice design parameters (specifically channel width and conductive layer thickness) are optimized through 3D-CFS to maximize sensing performance.
@@ II Principle of EIT-based tactile sensing
The principle of EIT-based tactile sensing is illustrated in Fig. 1. When localized pressure is applied to a sample substrate, it induces changes in conductivity within the touched regions, which in turn alter the electric potentials measured at the boundary electrodes. The objective of EIT-reconstruction in tactile sensing is to estimate the pressure-induced conductivity changes $\boldsymbol{\sigma}$ in the sensing region $\boldsymbol{\Omega}$ based on the voltage measurements $\boldsymbol{V}$. EIT involves solving both a forward and an inverse problem. The forward problem predicts surface potentials for a given conductivity and applied current. This relationship is governed by:
@@
III Simulation-Guided Structural Optimization and Experimental Validation
Figure 2: Electric field distribution for (a) uniform conductivity layer and (b) lattice-patterned conductivity layer.
III-A Physical Mechanism
Compared to sensors with a uniformly conductive medium, lattice-pattern EIT tactile sensors exhibit more pronounced voltage changes. This enhanced sensitivity primarily arises from two interrelated factors: increased local current density and modified electric field distributions [23].
In a lattice-patterned sensor, the conducting material is restricted to predetermined channels arranged in a lattice grid pattern while the rest of the area is non-conductive. This design forces current to flow primarily through a limited number of channels, thereby increasing the local current density. When external pressure is applied, the geometry of these channels changes, altering their cross-sectional area and shape, thereby yielding significant local resistance changes.
The electric field lines in traditional EIT sensors are distributed relatively uniformly throughout the conducting medium (Fig. 2a). In contrast, lattice-patterned structures induce sharper electric field gradients within the conductive channels (Fig. 2b). Upon external pressure, the deformation of the channels further intensifies these localized field gradients, amplifying the sensor’s impedance response.
III-B 3D-CFS
Figure 5: Visualization of lattice pattern geometries under different conductive channel widths. (a) $w=2$ mm; (b) $w=4$ mm (optimized); (c) $w=6$ mm. Blue regions represent the conductive hydrogel channels, while white regions represent the silicone area.
@@ III-A Physical Mechanism
As shown in Fig. 3a, the simulated sensor model is designed to be 100 × 100 × 9 mm and consists of three layers and 16 evenly distributed boundary electrodes. The substrate and sealing layers are non-conductive, and the sealing layer has a fixed thickness of 2 mm. The thickness of the substrate layer is changed with the intermediate conductive lattice layer to maintain a constant total model thickness of 9 mm. To evaluate and optimise sensor sensitivity, six different lattice channel widths and five different conductivity layer thicknesses were simulated to assess the influence of lattice channel widths and conductive layer thickness on sensor sensitivity.
As shown in Fig. 3b, three different touch conditions are applied by varying the number of touch points from one to three. In each condition, 100 sets of touch phantoms are randomly generated, with press depths from 1 mm to 5 mm and touch radius from 6.25 mm to 8.75 mm.
@@
Different lattice parameters were calculated, with the conductive layer thickness from 1 mm to 5 mm and the channel width from 0 mm (no lattice pattern) to 6 mm. The result is as shown in Fig. 4, each average value represents the average relative voltage change from 300 random touch phantoms across three touch conditions. We defined the optimal structural parameters as a conductive layer thickness of $t=3$ mm and a conductive channel width of $w=4$ mm. These values were selected because they maximized the average relative voltage change $V^{\text{rel}}$ under touch indentation, while the other fabricated variants are considered “non-optimized” for comparison. To provide a more intuitive understanding of the geometric differences, we visualized the lattice patterns corresponding to the optimized configuration (Fig. 5b, $w=4$ mm) and two representative non-optimized variants (Fig. 5a, $w=2$ mm; Fig. 5c, $w=6$ mm).
III-C Sensor Fabrication
Figure 6: Hydrogel fabrication process. (a) The solution contained 21.5 weight percent (wt%) lithium chloride (Sigma-Aldrich), 8 wt% acrylamide (Sigma-Aldrich), 32.27 wt% deionised water and 37 wt% ethylene glycol (Sigma-Aldrich) was prepared. Ethylene glycol maintains hydrogel pliability across a wide temperature range, essential for stable sensor performance. (b) The solution was stirred at 60 $^{\circ}\mathrm{C}$ and 700 rpm for two hours to achieve uniform mixing and dissolution. (c) To initiate a polyacrylamide-based network, 1 wt% polyacrylamide (Sigma-Aldrich) was introduced, providing mechanical stability and enhancing sensor durability. (d) The solution was stirred overnight at 60 $^{\circ}\mathrm{C}$ and 700 rpm to ensure complete dissolution and formation of a uniform polymer network. (e) A crosslinking agent, 0.15 wt% N,N’-methylenebisacrylamide (Sigma-Aldrich), was added to control the mechanical strength of the gel. Additionally, 0.08 wt% 2-hydroxy-2-methylpropiophenone (Sigma-Aldrich), a photoinitiator, was added to enable UV curing of the hydrogel. (f) The mixture was shielded from light and stirred for an additional two hours at 60 $^{\circ}\mathrm{C}$ and 700 rpm to ensure complete homogenization.
@@ III-C Sensor Fabrication
Based on the optimal design parameters identified through the 3D-CFS simulations, a conductive layer thickness of 3 mm and a lattice channel width of 4 mm, we proceeded to fabricate the lattice-structured EIT tactile sensor accordingly. The sensing layer was fabricated from hydrogel, and the hydrogel precursor synthesis procedure was shown in Fig. 6. As shown in Fig. 7, the hydrogel and silicone layers were integrated to fabricate the complete tactile sensor (Fig. 7h).
To enhance durability and maintain long-term functionality, two mitigation strategies against water evaporation were adopted. First, lithium chloride was selected as the ionic conductor in the hydrogel precursor formulation due to its hygroscopic properties, which effectively suppress dehydration while maintaining high ionic conductivity. Second, the entire hydrogel layer was encapsulated with a stretchable silicone elastomer to serve as a physical barrier against moisture loss [24].
III-D Real-world Sensitivity Characterization and Validation
@@
To verify the optimized sensor parameters identified through CFS, we fabricated three sensors using the same procedures (refer to Fig. 7) and a consistent conductive layer thickness ($t=3$ mm): two lattice-patterned sensors with different lattice parameters ($w=4$ and $6$ mm) and one without the lattice structure for comparison. These sensors were subjected to a continuously increasing localized pressure with strains ranging from 0 mm to 4 mm, as illustrated in Fig. 8a. The pressure was applied through a circular indenter with a radius of 8.75 mm, simulating the typical contact area of a human fingertip. The pressing location, marked by the blue dot in the inset of Fig. 8a, remained consistent across all measurements. The results demonstrate that the sensor with optimized lattice parameters ($w=4$ mm) consistently shows superior sensitivity compared to the non-lattice structured sensor and lattice-structure sensor with non-optimal parameters ($w=6$ mm), aligning well with the simulation results. Specifically, at the press depth of 4 mm, the average relative voltage change $V^{\text{rel}}$ for the sensor with the optimized lattice ($w=4$ mm) reaches approximately 0.0692, compared to  0.0613 for the non-optimized lattice sensor ($w=6$ mm) and only  0.0375 for the non-lattice sensor. This corresponds to an 84.5% improvement over the non-lattice structure and a 12.9% enhancement compared to the non-optimized parameters ($w=6$ mm), clearly confirming the effectiveness of the optimized lattice pattern in enhancing sensitivity.
Additionally, tactile reconstruction results comparing the optimized lattice-patterned sensor and the sensor without the lattice channel are illustrated in Fig. 8b. The reconstructed images show that the optimized lattice-patterned sensor successfully reconstructs tactile information even at lower compressive strains, whereas the sensor without the lattice structure fails to identify tactile input at these shallow press depths. Furthermore, as the compressive strain increases, the lattice-patterned sensor consistently demonstrates superior reconstructions compared to the sensor without the lattice structure. These results further validate that the lattice-patterned sensor possesses higher sensitivity and enhanced tactile reconstruction performance.
III-E Durability Evaluation via Cycling Test
@@
IV Tactile reconstruction: Spatial Resolution Evaluation
Figure 9: Tactile reconstruction using experiment data. Images are normalized.
Figure 10: Tactile reconstruction of unseen real-world complex touches. Images are normalized.
@@ IV-A Dataset and Network Training
To generate the dataset for the lattice-structured EIT sensor, we constructed a finite element model (FEM) using COMSOL Multiphysics and Matlab. We divide the sensing region with a size of 100 $\times$ 100 mm2 into 2304 elements (48 $\times$ 48) and set the background conductivity $\boldsymbol{\sigma_{0}}$ to 0.00312 S/m. Tactile interactions were simulated using circular regions with diameters ranging from 7.5 mm to 27.5 mm, varying in both location and intensity. The conductivity within these touch regions is randomly assigned values between $0.05\times\sigma_{0}$ and $2\times\sigma_{0}$. The complete dataset comprises 50,000 samples, divided into five subsets containing 10,000 samples each. These subsets correspond to different numbers of touch regions, ranging from 1 to 5 touch areas per sample. Each subgroup is further partitioned into training, validation, and testing sets using a 7:2:1 ratio. To enrich the training dataset by supplementing uncollected location information, we used a data augmentation approach to convert a single EIT measurement into 8 distinct readouts [26]. To further improve generalisation, additive Gaussian noise with Signal-to-Noise Ratios (SNRs) of 35, 40, 45, 50, 55, 60 and 65 dB was added to the seven augmented measurements, respectively. Finally, this process yielded a total of 280,000 samples for training, 10,000 samples for validation, and 5,000 samples for testing.
@@ IV-B Reconstruction Results and Discussion
As shown in Fig. 9, from Phantom 1 to Phantom 8, we applied 1 to 5 contacts at different positions using a linear robotic platform (DLE-RG-0001). The reconstruction results demonstrate that the sensor with the DNN model successfully reconstructs various touch patterns with good accuracy. To quantitatively assess reconstruction quality, we evaluate the results using correlation coefficients (CC), peak signal-to-noise ratio (PSNR), relative image error (RE), and structural similarity indexes (SSIM) [27, 28]. Across all phantoms, the model achieves strong performance, with CC values reaching up to 0.9275, RE values as low as 0.3798, SSIM values up to 0.9660, and PSNR values reaching 29.0303 dB. These results indicate that the proposed sensor and model combination delivers high-fidelity reconstructions with strong consistency and visual clarity.
Beyond the above benchmark scenarios, we further validated the system’s performance on real-world measurements involving more complex and natural touch patterns. These include rectangular contact regions and annular touch profiles. Neither of them is included in the training dataset, so these cases serve as a strong test of the model’s generalization capability. As shown in Fig. 10, the DNN model successfully reconstructs these diverse tactile configurations with clear spatial patterns and consistent localization, demonstrating that the learned representation is robust to previously unseen contact geometries and pressure distributions. For instance, in Phantom 6 and Phantom 7, which correspond to rectangular and annular touch profiles, respectively, the model approximates these shapes using combinations of circular blobs. This behavior arises from the fact that the training data only included circular touch patterns; nonetheless, the model is still able to capture the overall geometry and spatial distribution.
This generalization stems from the training dataset incorporates a wide range of touch sizes, intensities, and noise perturbations, enabling the model to learn transferable spatial priors. Supplementary Video 1 further demonstrates the model’s real-time tactile reconstruction capabilities in these challenging scenarios, reinforcing its suitability for practical, interactive applications with arbitrary and dynamic touch inputs.
V Human–Machine Interaction Demonstration